OpenStudy (anonymous):

(4 pt) Write the inverse variation equation for the relationship: y varies inversely with x and y = 4 when x = 2.

OpenStudy (anonymous):

a: y= 8/x

OpenStudy (anonymous):

b: y=2x

OpenStudy (anonymous):

c: y=x+2

OpenStudy (anonymous):

d: y=1/2 x

OpenStudy (anonymous):

hello? please i need help